# Label-Guided In-Context Learning for Named Entity Recognition

### v1

@@ 5 Related Work
An emerging capability of LLMs is their ability to reason (Wei et al., 2023) and self-reflect (Asai et al., 2023) on their predictions, enabling them to correct errors. Recent research has explored this capability across a wide range of tasks, like code generation (Madaan et al., 2023) and planning (Shinn et al., 2023). Similar approaches have also been applied to NER (Wang et al., 2023a; Chen et al., 2024). These works primarily rely on LLMs’ internal knowledge for decision-making. In contrast, our approach grounds model predictions to the label-based training statistics. This grounding enables the model to capture fine-grained dataset-specific artifacts, akin to supervised methods, and helps close the performance gap between prompting-based and supervised approaches.
@@ 6 Conclusion
We present Deer, an ICL-based NER method that enhances demonstration selection through token-level, label-guided retrieval and improves prediction through targeted error reflection. Evaluated on five NER datasets with four LLMs, Deer consistently outperforms existing ICL methods and achieves performance comparable to supervised fine-tuning. Further analyses demonstrate its effectiveness on both seen and unseen entities and its robustness in low-resource settings.
@@ A.1 Deer Prompts
Training Data Stats: 0 entity, 44 context, 20 regular

### v2

@@ 5 Related Work
An emerging capability of LLMs is their ability to reason (Wei et al., 2023) and self-reflect (Asai et al., 2023) on their predictions, enabling them to correct errors. Recent research has explored this capability across a wide range of tasks, like code generation (Madaan et al., 2023) and planning (Shinn et al., 2023). Similar approaches have also been applied to NER (Wang et al., 2023a; Chen et al., 2024). These works primarily rely on LLMs’ internal knowledge for decision-making. In contrast, our approach grounds model predictions to the label-grounded training statistics. This grounding enables the model to capture fine-grained dataset-specific artifacts, akin to supervised methods, and helps close the performance gap between prompting-based and supervised approaches.
@@ 6 Conclusion
We present Deer, a training-free ICL method designed to help LLMs make more informed entity predictions. Deer leverages label statistics to identify tokens that are most informative for entity recognition, thereby improving demonstration retrieval and enabling effective error reflection. Evaluated on five NER datasets across four LLMs, Deer consistently outperforms existing ICL approaches and achieves performance comparable to supervised fine-tuning. Further analyses highlight the limitations of commonly used sentence embedding-based retrievers and demonstrate Deer’s effectiveness on both seen and unseen entities, as well as its robustness in low-resource settings.
@@ A.1 Deer Prompts
Training Data Stats: 0 entity, $4$ context, 20 regular
@@ A.2 Statistical Test on Sampled Datasets
Method BERTbase Qwen2.57B GPT-4o-mini GPT-4o
Fine-tune 84.0 - - -
Kate - 66.9 76.0 80.0
GPT-NER - 71.2 75.6 82.7
Deer (w/o ER) - 71.7 77.6 82.7
Table 14: Comparing our Deer with GPT-NER. Deer consistently outperforms GPT-NER across different LLMs. Notably, GPT-NER even underperforms its original BERT-based retriever, highlighting its limitations and impracticality.